OpenStudy (anonymous):

What is the fifth term of the geometric sequence? a1 = 120, a2 = 36, a3 = 10.8, a6 = 0.2916

OpenStudy (anonymous):

a2/a1=a3/a2=a5/a6

OpenStudy (anonymous):

@HARSH123 explain please

OpenStudy (anonymous):

.3=.3=.2916/a5

OpenStudy (anonymous):

solve this